UT Nov 19 2021 - D/E 1st half, LBTI Science 2nd half



Night Info

  • Night Center 00:05 - handover to LBTI 23:53hr
  • Sunset 17:22
  • Twilight Ends 18:42
  • Twilight Starts 05:28
  • Sunrise 06:48
  • Full Moon

Night Plan

  • Starting config bino-MODS
  • SpecStd Feige110 (Dual Grating) in twilight
  • M31 Nova 18:42-20:42hr
    • File Location: /home/lbto/Service_Observing/2021B_D_E/MODS/RMWagner_M31_Nova
    • Use either m31n2008_pa0_1pt0_Late.acq or m31n2008_pa0_1pt2_Late.acq - 120sec acquisition and 1" or 1.2" slit depending on seeing
    • User either m31n2008_1pt0_extended.obs or m31n2008_1pt2_extended.obs for science portion - 5x1200sec exposures
  • Reconfigure for LBCs 20:45-21:10hr
  • LBC Timing Tests 21:10-23:53hr
  • Reconfigure for LBTI 23:3
  • Poor Conditions Backup:
    • Bino-LUCI
    • Local Mergers
      • OT File: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/COBandhead_Mergers_V29.xml
      • Any conditions (no priority order)
      • NGC 6598 - 1hr bino (repeat script 1x) if really bad, go longer - twilight thru 20:30hr
      • NGC 7585 - 1hr bino (repeat script 1x) if really bad, go longer - twilight until 23hr
      • UGC 11905 - 1hr bino (repeat script 1x) if really bad, go longer - twilight until 23hr
      • Arp 230 - 1hr bino (repeat script 1x) if really bad, go longer - 18:30hr thru handover
        • telluric rises 19:30hr thru handover

Overview

MODS --- Feige 110 and M31 Nova

22:40 MODS brough up and awake, put in observing mode for Josh's test preset. Since this is an eng night I thought I would be daring and use MODS on OBS5.
  • Requires export PATH=/lbt/lbto/mods/bin:$PATH in terminal, but other than that all appears to be coming up ok. This was similar to ROBS in its test phase.
22:55 Running MODS simSnap and turning on LBC's. Red channel pressures lost so lbckill/lbcstart and powered up again.

23:02 Taking 25 LBC Bias, and MODS 8K Bias

modsDisp error on OBS5:
Traceback (most recent call last):
  File "/lbt/lbto/mods/bin/modsDisp", line 398, in <module>
   watcherLoop()
  File "/lbt/lbto/mods/bin/modsDisp", line 312, in watcherLoop
    md.set('scale linear')
  File "/usr/lib64/python3.7/site-packages/pyds9/pyds9.py", line 646, in set
    self._selftest()
  File "/usr/lib64/python3.7/site-packages/pyds9/pyds9.py", line 571, in _selftest
    1):
  File "/usr/lib64/python3.7/site-packages/pyds9/xpa.py", line 209, in xpaaccess
    raise ValueError(errmsg)
ValueError: XPA$ERROR: no response from server during handshake (?:?)

Completed dual grating arcs, dual grating 1" flats, dual grating 1.2" flats

0:40 Josh is executing pointing and collimation near Feige 110, still a bit bright for collimation but pointing all good. The clouds are starting to come in.

0:52 Tryin collimation preset again. Seeing is ~1-1.1" on the guiders.

0:54 Sending preset for Feige 110.

1:05 Aligned and waiting for 12 degree twilight to start science. There was an issue with modsAlign on obs5. It was unable to find images. Running modsAlign on obs2.

MODS1 commanded offsets for Acq: -0.228 12.184, additional correction -0.437 -0.123
MODS2 commanded offsets for Acq: 4.263 9.298, additional correction -0.413 -0.059

1:12 Science script sent of Feige 110. Seeing 0.8" on the guiders, 0.76" on the DIMM. 1:13 12 degree twilight

mods[1/2]r.20211119.0026-28.fits
mods[1/2]b.20211119.0023-25.fits

Clouds have temporarily cleared, some thin cirrus

1:20 Seeing is bouncing between 0.8-0.95" on the DIMM and guiders. Clouds along the western and northern horizon looming. Will send the 1" longslit script for M31 Nova.

1:26 Sending preset to m31 Nova m31n2008_pa0_1pt0_Late_v2.acq. Exposure time bumped to 150sec to account for the clouds starting to form in the north. Seeing 0.93" on the DIMM, 0.78" on the guiders. Clouds have moved in directly covering this current pointing. Target barely visible, particularly in MODS1, clouds were just thick during acq exposures but appear to once again be thinning out.. Increased Exp time to 180sec for mods1

MODS1 commanded offsets for Acq:- -0.118, 10.415
MODS2 commanded offsets for Acq: 4.312, 8.360

1:42 18 degree twilight

1:46 Starting science for m31 nova, 4x1800sec exposures. Clouds appear to be thinning out again with only thin currus overhead. Seeing 0.97" on the DIMM and 0.6-0.7" on the guider.

mods[1/2]r.20211119.0032.fits
mods[1/2]b.20211119.0026.fits

Cirrus wafting through in the north periodically, but no significant drops in flux (nothing like what was seen during acquisition).

2:19 First images read out. H alpha clearly visible in the red images. mods_quick reduce issue on obs2&3 still present (IT 8027). MODS_QUICKREDUCE: Invalid FITS Header for file mods1r.20211119.0032.fits

3:23 Starting the 4th and final exposure, that moon is bright! Some thin cirrus, less than 0.5 mag extinction. Seeing has been pretty stable ~0.6-0.8".

3:40 More cirrus moving in, its starting to get a bit thicker. Seeing 0.6-0.7" on the guiders

3:57 Reconfiguring to LBC's. Hatch closed on the MODS. Heejoo is starting TMS lasers

LBCs

Timing Tests

04:10 Copied reference from 18 Nov 2021 to /tmp

[telescope@robs 20211118]$ cp telescope_left_tmsoffset.dat /tmp
[telescope@robs 20211118]$ cp telescope_right_tmsoffset.dat /tmp
[telescope@robs 20211118]$ cp tms_dx_ref.dat /tmp
[telescope@robs 20211118]$ cp tms_sx_ref.dat /tmp

04:14 Slewing to copointing field for G30 04:25 and wait for group rotation updates.

There was an LBC Blue controller error (bad input voltage on line 9). It did not automatically recover and I reinitialized the controller by powering off/on the controller through power.py and then running lbckill/lbcstart.

04:21 pause TMS and run dofpia.

04:45 Slewing to copointing field for G30 04:45 ... I tried a couple different fields for FPIA that were close to the satellite coordinates, but with the moon and cirrus, no pupils were found above the background.

Finally went to a copointing/collimation field for one of the PHAs with lots of bright stars

04:50 First corrections:

  -532    -93   -258    274    265   -148    -22   1.02   0.31    blue
  -173     12    209    -14    110    -81     16   0.62   0.14     red

Blue camera:

There is a 35-sec delay between resume & the start of exposure

04:57 Slewing to G19 05:00. The satellite is on chip 2

05:05 on chip 2

Assume 38 seconds( - this time 3 sec early)

05:10 on chip2

05:15 on chip 2 (time very close to 0515)

05:20 on chip2

Red camera:

There is an ~30-sec delay between resume and the start of the exposure (assumed first a 20 sec delay)

05:25 --- 20 sec is too short, object near top of chip 3. delay was 32

05:30 on chip 2

05:35 on chip 2

05:40 on chip 2 maybe 27 instead of 30 sec delay for next time

05:45 on chip 2

Repeat TMS testing using the PHAs that were observed on 20211115 UT and 20211117 UT

2004 QB (elevation 39 degrees)

05:58 Slewing to copointing field for 2004 QB to catch it as it's rising
Radial star offsets from rotator centers: BLUE 27.5" and RED 44.0"

lbcrangebal:
-------------------------------------------------------------------------
COPOINTING: B=60636 R=60604
Pointing updates: delta_IE = -11.67", delta_CA = 6.00"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.52     1.54    33.02   -11.26
              DX:   -0.51    -1.50   -32.11    10.86
-------------------------------------------------------------------------

06:13 Starting the observation 2004 QB: At only 38 deg from the moon, the exposure time of 60-s was too much. Dropped it to 30 and then switched filters from g+r to r+i and dropped the time to 18 sec.

2013 MW6 (elevation 82 deg)

06:37 Slewed to a high elevation field (82 deg) (03:41, 29:05)
  • 3 pairs of images here.
  • object name is 2021 UQ1 but the target is actually 2013 MW6

2021 UQ1 (elevation 57 deg)

06:43 Now going to 2021 UQ1

064524 and 064514 were taken before the mirror settled. I took one more pair of exposures before we had to quit. The red IQ was not that good.

06:53 ran fpia (but the control C did not take so do not trust the corrections based on the 2nd and 3rd pupil images).

 94   -186   -209    341    390   -103     24    y  1.35  0.27                 /lbt/data/new/lbcb.20211119.065021.fits  V-BESSEL
-3146    216   -210    451   -246     93     24    y  0.91  1.27                 /lbt/data/new/lbcr.20211119.065017.fits  r-SLOAN 

SX Back-face corrections:    -777.00 Z4 (nm) or  -0.029448 dZ (mm) from 4.8445 to 6.8564 hours UT
DX Back-face corrections:    -939.90 Z4 (nm) or  -0.035622 dZ (mm) from 4.8450 to 6.8569 hours UT

Summary

With the LBCs, we obtained five 0.5-sec exposures with LBCB and then with LBCR of the satellite G19 in order to check the accuracy of the timestamp that is written to the LBC headers as date_obs. We used the TMS reference from last night, 20211118 UT. On DX, all Zernikes were within the convergence criteria, but on SX, Z4 was large, -532 nm. We did not run lbcrangebal, but it was apparent from the pupil images that at least the copointing was off.

We spent the remaining time running TMS in active mode and observing the 3 PHAs that had been observed on 15 & 17 Nov UT. We copointed on the first field, of 2004 QB, at 38 deg elevation, and then observed it for about 20 minutes before moving to 2013 MW6 (82 deg elevation) and finally to 2021 UQ1 (57 deg) for which we got only two pairs of images. The first pair were taken before the mirrors settled. Immediately after taking the second pair of images we ran FPIA. Unfortunately, the active loop was not paused (the control-C did not "take"), but the Zernikes computed using the first pair of pupils should be fine and, while for LBCB, these were all relatively small, for LBCR, there was a large Z4 correction indicated (also a relatively large Z7 one).

-- BarryRothberg - 18 Nov 2021
Topic revision: r9 - 21 Nov 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback